# End-to-End Large Portfolio Optimization for Variance Minimization with Neural Networks through Covariance Cleaning

### v1

@@ 3.1 Lag-Transformation module
It is reasonable to assume that past observations do not influence the future risk uniformly. To make our NN lag aware, we design a specific module that, before the input returns reach the subsequent blocks, they are processed through a learnable lag-transformation system that assigns a distinct transformation to every look-back lag. Classical risk models define a fixed exponential kernel [46, 20]; our formulation generalises that kernel by letting the NN determine its shape by training on the real data.
@@
with $\bm{\alpha}=\{\alpha_{t}\}_{t=1}^{\Delta t_{\text{in}}}$ and $\bm{\beta}=\{\beta_{t}\}_{t=1}^{\Delta t_{\text{in}}}$ learned by back-propagation. The coefficient $\alpha_{t}$ rescales observations at lag $t$. The coefficient $\beta_{t}$ modulates their outlier sensitivity by controlling the saturation of the hyperbolic tangent. When $\beta_{t}$ is small the mapping is nearly linear ($\tanh(x)\approx x$ if $x$ is small) and the return enters almost unchanged; when $\beta_{t}$ is large the mapping approaches a symmetric step that clips extreme values at ±$\pm\alpha_{t}/\beta_{t}$. In this way, the module can emphasise or attenuate individual lags and dampen outliers simultaneously.
The factor $252$ annualises the input and stabilises gradient computations under single-precision arithmetic. Because $\alpha_{t}$ and $\beta_{t}$ are updated jointly with all other network parameters, the module can learn any monotone or non-monotone lag profile, with the classical exponential decay arising as a special case. The output transformed matrix $\widetilde{\bm{R}}$ serves as the input to all subsequent modules.
@@ 3.2 Eigenvalue Cleaning with Neural Networks
To design an architecture that more effectively mitigates the curse of dimensionality, we exploit the exact mathematical mapping provided by the Coulomb-gas formulation of RMT [34], which models sample eigenvalues as one-dimensional charges under three forces: mutual logarithmic repulsion, confinement by the population spectrum, and hard-edge repulsion at zero. The confinement potential reflects the full population spectrum: when eigenvalues are well separated relative to the effective sample length $\Delta t_{\mathrm{in}}$, it is dominated by the nearest population eigenvalue; otherwise, it arises from a complex superposition of bulk contributions. The mutual repulsion between sample eigenvalues scales as $|\widetilde{\lambda}_{i}-\widetilde{\lambda}_{j}|^{-1}$ and therefore decays with rank separation, while the hard-edge term enforces nonnegativity but weakens as $\Delta t_{\mathrm{in}}$ approaches $n$. Consequently, the shift of each $\widetilde{\lambda}_{i}$ away from its population counterpart is governed primarily by interactions with the sample eigenvalues that lie within its local spectral vicinity, a region whose extent depends on the local spacing around $\widetilde{\lambda}_{i}$, while the cumulative influence of all more distant eigenvalues can be approximated as a coarse-grained contribution. Although the resulting spectral density and eigenvector-overlap dynamics could in principle be derived by solving the coupled Hamiltonian equations associated with this analogy, we instead incorporate only the implied locality as an inspiration for a structural prior in our neural architecture, which should lie, in fact, between the extremes of Deep Sets and attention architectures.
@@
$g^{\rightarrow}(\bm{h}^{\rightarrow}_{i-1},\bm{m}^{\rightarrow}_{i-1},\widetilde{\lambda}_{i},q)$ =$=\{\bm{h}^{\rightarrow}_{i},\bm{m}^{\rightarrow}_{i}\},$ $\text{with}\quad i=1,\ldots,n,$ (14)
$g^{\leftarrow}(\bm{h}^{\leftarrow}_{i+1},\bm{m}^{\leftarrow}_{i+1},\widetilde{\lambda}_{i},q)$ =$=\{\bm{h}^{\leftarrow}_{i},\bm{m}^{\leftarrow}_{i}\},$ $\text{with}\quad i=n,\ldots,1,$ (15)
where $\bm{h}^{\rightleftharpoons}_{i}=\left\{h^{\rightleftharpoons}_{i,1},\dots,h^{\rightleftharpoons}_{i,k}\right\}$ and $\bm{m}^{\rightleftharpoons}_{i}=\left\{m^{\rightleftharpoons}_{i,1},\dots,m^{\rightleftharpoons}_{i,k}\right\}$ are the hidden and cell states. Setting the input states to zero at the start of the sequences, Eq. (14) is applied from the smallest to the largest eigenvalue, whereas Eq. (15) proceeds in the opposite direction.
@@
After scanning the spectrum, the BiLSTM outputs an $n\times 2k$ matrix $\bm{H}$ whose $i$-th row $\bm{h}_{i}$ concatenates column-wise the forward $\bm{h}^{\rightarrow}_{i}$ and backward $\bm{h}^{\leftarrow}_{i}$ hidden vectors. A dense layer shared across ranks projects $\bm{h}_{i}$ onto the positive real line through a softplus activation
$\lambda_{i,\text{NN}}^{-1}=\operatorname{softplus}(\mathbf{w}^{\top}\bm{h}_{i}+b),\qquad i=1,\dots,n,$ (16)
with trainable parameters $\mathbf{w}\in\mathbb{R}^{2k}$ and $b\in\mathbb{R}$. The softplus retains differentiability at the origin and avoids hard saturation, ensuring stable gradients even for very small eigenvalues. To remove the arbitrary overall scale, we normalise such that $\sum_{i=1}^{n}\lambda_{i,\text{NN}}^{-1}=n$. Portfolio weights computed from a RIE do not change if all the eigenvalues are multiplied by the same constant, so this normalization places every instance on the same scale and allows direct comparison in downstream statistical tests without altering the resulting portfolio weights.
In summary, the module comprises a total of $4338k^{2}+26k+1=34,$ 433 learnable parameters ($k=64$). Both LSTM cell contributes $8k^{2}+24k$ parameters, and the final dense layer adds $12k+$ 1. If model compactness is a priority, the LSTM units may be replaced by two Gated Recurrent Units (GRUs). Standard LSTMs employ separate memory and hidden channels, $\bm{m}$ and $\bm{h}$, respectively, to mitigate vanishing and exploding gradients. The cell state $\bm{m}$, modulated by input, forget, and output gates, can preserve long-range information almost unchanged over many time steps, while the hidden state $\bm{h}$ undergoes more aggressive step-wise updates and serves as the immediate output. This arrangement improves numerical stability and captures both global and local spectral structure. That said, this dual-channel design is not strictly necessary: in our experiments, a simplified GRU, which merges long and short-term memory into a single $\bm{h}$ performed comparably well. The GRU reorganizes its gating mechanism so that the hidden state alone can reliably encode long-range dependencies, offering a lighter yet effective alternative. Substituting LSTM cells with GRUs reduces the total parameter count to $16k^{2}+19k+$ 1.
@@ 3.3 Multilayer Perceptron for Inverse Marginal Volatilities
Each standard deviation is then passed through an MLP network with $3$ hidden layers of $(64,32,16)$ neurons and a Leaky-ReLU activation function. The output of the network is a single neuron with a softplus activation function, which guarantees that the output is positive. In total, this module comprises $2,753$ learnable parameters. The output of this branch is then standardized by dividing it by the average inverse output of the network, i.e.,
$\sigma_{i,\text{NN}}^{-1}\rightarrow\frac{\sigma_{i,\text{NN}}^{-1}}{\frac{1}{n}\sum_{j=1}^{n}\sigma_{j,\text{NN}}^{-1}}.$ (18)
@@ 4.1 Selection of the Investable Universe
We describe here the procedure used to select assets for investment from the universe of US equities that were listed on the NYSE or NASDAQ at any point from 1 January 1990 to 31 December 2024. This universe, therefore, includes American Depositary Receipts (ADRs) and explicitly excludes funds and exchange-traded funds. To avoid data leakage, for each trading day $t$, we rely only on information available up to $t-1$ when determining the eligible stocks. We further exclude any security for which a planned delisting occurs within the subsequent $\Delta t_{\text{out}}=5$ trading days.
@@
On the recent trading history, we require that 100% of the auctions are executed and at least 1% of the shares outstanding and 1% of the market capitalisation are traded on average over the last five days. This ensures that the stock is liquid enough to be traded with a low impact on the market. The day before the trade, we check that the number of shares outstanding is greater than 5M, to avoid micro-shares. Finally, on the day before trading, we require that each stock’s price lies between 10 \$ and 2,000 \$. Stocks trading at relatively low price levels incur high per-share transaction costs, whereas those at relatively high price levels often exhibit reduced trading activity, impeding efficient execution.
@@
Finally, to reduce redundancy, we keep only one class of shares for each company, the one with the largest market capitalization on the previous day. A similar filter is executed for those stocks with an in-sample correlation greater than 0.95 [55]. The final selection is done by ranking the stocks by their market capitalization and selecting the top $n=1,000$ stocks for every trading day $t$. The final panel, spanning January 1990 to December 2024, contains roughly $2.7$ millions of stock–day pairs and $3,837$ unique stocks.
@@ 4.2 Training Procedure
The proposed architecture is designed to be independent of the number of assets $n$. This means that the same model can be applied to different subsets of assets without the need for retraining. However, we know from RMT that the filtering should depend on the aspect ratio $q=n/\Delta t_{\text{in}}$; therefore, during the training phase, the NN should be fed with examples of different $n$ if we aim to obtain a model that generalizes well to larger portfolio sizes. This is done by randomly and uniformly sampling the number of assets $n$ from a $[50,350]$ range on every batch of the training. The number of days $\Delta t_{\text{in}}$ is fixed to $1200$ trading days, in fact, the presence of the module of Sec. 3.1 includes $2\Delta t_{\text{in}}$ trainable variables $(\alpha_{t},\beta_{t})$, which are associated to specific past lags. In the future version, this could be generalized to a variable number of past lags, but it would require a more complex architecture.
@@
According to the training procedure, the probability that NN will be trained on the same dataset is very low for $n$ sufficiently smaller than the investable universe on a given day. This allows us to train the NN without the need to monitor the performance on a validation set for early stopping. However, to further reduce the risk of overfitting, we used a decay factor of the learning rate, which is set to $10^{-4}$ and drops every batch by a factor of $0.99^{1/500}$, and we clip the gradients by norm to $1.0$. The gradient is then optimized with the Adam optimizer with an automatic rescaling of the loss function. We trained all our models with $100$ epochs, with $500$ steps per epoch, and a batch size of $32$ samples. To ensure that the convergence is stable, for each of the 24 years, we trained our NN 10 times independently, so finally, we obtained 240 trained NNs.
@@
For further evidence of robustness against overfitting, Fig. 2 displays the evolution of training and validation losses over successive epochs. The loss trajectories remain consistent across independent runs on the same data interval. Both training and validation losses decrease monotonically. Monotonic decline of the training loss is expected; the validation loss, by contrast, often increases once overfitting sets in. After $60$ epochs, the training loss plateaus, exhibiting an average change of less than $3.5\%$ over the final forty epochs. The validation loss continues to decline at a faster pace, falling by approximately $6\%$ during the same period. These results indicate that terminating training around $100$ epochs would probably incur minimal overfitting risk.
@@ 4.3.1 Lag-Transformed Returns
After the calibration, we can analyze the learned weighting factors $\alpha_{t}$ and $\beta_{t}$ of Eq. (13) as a function of the lag $t$ to understand how the NN is weighting and transforming the past returns. Given that we ran $10$ independent training runs for each of the $24$ years of the validations, we obtained $240$ trained models. We can then compute the average of the weighting factors over the trained models, which are shown in Fig. 3.
@@
Interestingly, the rescaling factor $\alpha_{t}$ seems to follow a power law with an exponent of approximately $-0.2$, which in substance means that the most recent past is weighted slightly more than the oldest past. In fact, half of the contribution is given from the $40\%$ most recent past. The emergence of the power law is quite surprising, being the most widely used decay factor based on Exponentially Weighted Moving Average (EWMA) [46, 20], whereas the power law decay would instead imply a Hyperbolic Weighted Moving Average (HWMA). A model based on HWMA was proposed in Ref. [56], but to our knowledge, it is not very popular yet.
The clipping factor $\beta_{t}$ shows values within the range $[0.4,1.0]$. Interestingly, the clipping factor $\beta_{t}$ seems to saturate for larger lags. To have an idea about the range of clipping, for the farthest data points, assuming $\alpha_{t}=1$, a daily return of $1\%$ is clipped to approximately $0.4\%$ while the same return in the most recent past is clipped to approximately $0.9\%$.
@@ 4.3.2 Eigenvalue Sensitivity Analysis
The surprising result of the AO estimator [30, 29] is that a set of fixed eigenvalues can be used to obtain an estimator for the correlation matrix that outperforms all the RMT-based estimators for portfolio optimization. Moreover, such an estimator does not depend on the assets used in the calibration or the exchange, country, or time period. The authors of Ref. [30] hypothesized that the non-stationarity of the system is so strong that, unless eigenvectors are corrected, very little information can be extracted from the sample eigenvalues. In our case, we used an NN trained to attempt extracting specifically the non-stationary relationship between $\widehat{\bm{C}}_{\text{in}}$ and $\bm{C}_{\text{out}}$. If the above statement is true, and the NN does not overfit the calibration data, we should expect an output that is independent of the assets in the input.
Figure 4: Eigenvalue sensitivity analysis. Left: median of the eigenvalues as a function of the rank, the colored bands indicate a $95\%$ percentile coverage. Right: Standard deviation of the log eigenvalues as a function of the rank. The eigenvalues are computed on 1024 samples from the full calibration dataset.

### v2

@@ 3.1 Lag-Transformation Module
It is reasonable to assume that past observations do not influence the future risk uniformly. To make our NN lag aware, we design a specific module that, before the input returns reach the subsequent blocks, processes them through a learnable lag-transformation system that assigns a distinct transformation to every look-back lag. Classical risk models define a fixed exponential kernel [55, 18]; our formulation generalizes that kernel by letting the NN determine its shape by training on the real data.
@@
with $\bm{\alpha}=\{\alpha_{t}\}_{t=1}^{\Delta t_{\text{in}}}$ and $\bm{\beta}=\{\beta_{t}\}_{t=1}^{\Delta t_{\text{in}}}$ learned by back-propagation. The coefficient $\alpha_{t}$ rescales observations at lag $t$. The coefficient $\beta_{t}$ modulates their outlier sensitivity by controlling the saturation of the hyperbolic tangent. When $\beta_{t}$ is small, the mapping is nearly linear ($\tanh(x)\approx x$ if $x$ is small) and the return enters almost unchanged; when $\beta_{t}$ is large, the mapping approaches a symmetric step that clips extreme values at ±$\pm\alpha_{t}/\beta_{t}$. In this way, the module can emphasise or attenuate individual lags and dampen outliers simultaneously.
The factor 252252252 annualises the input and stabilises gradient computations under single-precision arithmetic. Because $\alpha_{t}$ and $\beta_{t}$ are updated jointly with all other network parameters, the module can learn any monotone or non-monotone lag profile, with the classical exponential decay arising as a special case. The output transformed matrix $\widetilde{\bm{R}}$ serves as the input to all subsequent modules.
@@ 3.2 Eigenvalue Cleaning with Neural Networks
To design an architecture that more effectively mitigates the curse of dimensionality, we exploit the exact mathematical mapping provided by the Coulomb-gas formulation of RMT [32], which models sample eigenvalues as one-dimensional charges under three forces: mutual logarithmic repulsion, confinement by the population spectrum, and hard-edge repulsion at zero. The confinement potential reflects the full population spectrum: when eigenvalues are well separated relative to the effective sample length $\Delta t_{\mathrm{in}}$, it is dominated by the nearest population eigenvalue; otherwise, it arises from a complex superposition of bulk contributions. The mutual repulsion between sample eigenvalues scales as $|\widetilde{\lambda}_{k}-\widetilde{\lambda}_{l}|^{-1}$ and therefore decays with rank separation, while the hard-edge term enforces nonnegativity but weakens as $\Delta t_{\mathrm{in}}$ approaches $n$. Consequently, the shift of each $\widetilde{\lambda}_{k}$ away from its population counterpart is governed primarily by interactions with the sample eigenvalues that lie within its local spectral vicinity, a region whose extent depends on the local spacing around $\widetilde{\lambda}_{k}$, while the cumulative influence of all more distant eigenvalues can be approximated as a coarse-grained contribution. Although the resulting spectral density and eigenvector-overlap dynamics could in principle be derived by solving the coupled Hamiltonian equations associated with this analogy, we instead incorporate only the implied locality as an inspiration for a structural prior in our neural architecture, which should lie, in fact, between the extremes of Deep Sets and attention architectures.
@@
$\bm{g}^{\rightarrow}(\bm{h}^{\rightarrow}_{k-1},\bm{m}^{\rightarrow}_{k-1},\widetilde{\lambda}_{k},q)$ =$=\{\bm{h}^{\rightarrow}_{k},\bm{m}^{\rightarrow}_{k}\},$ $\text{with}\quad k=1,\ldots,n,$ (15)
$\bm{g}^{\leftarrow}(\bm{h}^{\leftarrow}_{k+1},\bm{m}^{\leftarrow}_{k+1},\widetilde{\lambda}_{k},q)$ =$=\{\bm{h}^{\leftarrow}_{k},\bm{m}^{\leftarrow}_{k}\},$ $\text{with}\quad k=n,\ldots,1,$ (16)
where $\bm{h}^{\rightleftharpoons}_{k}=\left\{h^{\rightleftharpoons}_{k,1},\dots,h^{\rightleftharpoons}_{k,\omega}\right\}$ and $\bm{m}^{\rightleftharpoons}_{k}=\left\{m^{\rightleftharpoons}_{k,1},\dots,m^{\rightleftharpoons}_{k,\omega}\right\}$ are the hidden and cell states. Setting the input states to zero at the start of the sequences, Eq. (15) is applied from the smallest to the largest eigenvalue, whereas Eq. (16) proceeds in the opposite direction.
@@
After scanning the spectrum, the BiLSTM outputs an $n\times 2\omega$ matrix $\bm{H}$ whose $k$-th row $\bm{h}_{k}$ concatenates column-wise the forward $\bm{h}^{\rightarrow}_{k}$ and backward $\bm{h}^{\leftarrow}_{k}$ hidden vectors. A dense layer shared across ranks projects $\bm{h}_{k}$ onto the positive real line through a softplus activation
$\lambda_{k,\text{NN}}^{-1}=\operatorname{softplus}(\mathbf{a}^{\top}\bm{h}_{k}+b),\qquad k=1,\dots,n,$ (17)
with trainable parameters $\mathbf{a}\in\mathbb{R}^{2\omega}$ and $b\in\mathbb{R}$. The softplus retains differentiability at the origin and avoids hard saturation, ensuring stable gradients even for very small eigenvalues. To remove the arbitrary overall scale, we normalise such that $\sum_{k=1}^{n}\lambda_{k,\text{NN}}^{-1}=n$. Portfolio weights computed from a RIE do not change if all the eigenvalues are multiplied by the same constant, so this normalization places every instance on the same scale and allows direct comparison in downstream statistical tests without altering the resulting portfolio weights.
In summary, the module comprises a total of $8\omega^{2}+26\omega+1=34,433$ learnable parameters ($\omega=64$). Both LSTM cell contributes $8\omega^{2}+24\omega$ parameters, and the final dense layer adds $2\omega+1$. If model compactness is a priority, the LSTM units may be replaced by two Gated Recurrent Units (GRUs). Standard LSTMs employ separate memory and hidden channels, $\bm{m}$ and $\bm{h}$, respectively, to mitigate vanishing and exploding gradients. The cell state $\bm{m}$, modulated by input, forget, and output gates, can preserve long-range information almost unchanged over many time steps, while the hidden state $\bm{h}$ undergoes more aggressive step-wise updates and serves as the immediate output. This arrangement improves numerical stability and captures both global and local spectral structure. That said, this dual-channel design is not strictly necessary: in our experiments, a simplified GRU, which merges long and short-term memory into a single $\bm{h}$ performed comparably well. The GRU reorganizes its gating mechanism so that the hidden state alone can reliably encode long-range dependencies, offering a lighter yet effective alternative. Substituting LSTM cells with GRUs reduces the total parameter count to $6\omega^{2}+19\omega+1$.
@@ 3.3 Multilayer Perceptron for Inverse Marginal Volatilities
Each standard deviation is then passed through an MLP network with 333 hidden layers of (64,32,16)(64,32,16)( 64 , 32 , 16 ) neurons and a Leaky-ReLU activation function. The output of the network is a single neuron with a softplus activation function, which guarantees that the output is positive. In total, this module comprises 2,7532,7532 , 753 learnable parameters. It is important to stress that we process each asset’s estimated volatility independently through the same MLP, taking one scalar at a time rather than feeding the entire vector of $n$ volatilities jointly. This ensures that the network’s parameters remain fixed and do not scale with the number of assets, while still learning a universal transform applicable to every marginal volatility. The output of this branch is then standardized by dividing it by the average inverse output of the network, i.e.,
$\sigma_{i,\text{NN}}^{-1}\mapsto\frac{\sigma_{i,\text{NN}}^{-1}}{\frac{1}{n}\sum_{j=1}^{n}\sigma_{j,\text{NN}}^{-1}}.$ (19)
@@ 4.1 Selection of the Investable Universe
We describe here the procedure used to select assets for investment from the universe of US equities that were listed on the NYSE or NASDAQ at any point from 1 January 1990 to 31 December 2024. This universe, therefore, includes American Depositary Receipts (ADRs) and explicitly excludes funds and exchange-traded funds. To avoid data leakage, for each trading day $t$, we rely only on information available up to $t-1$ when determining the eligible stocks. We further exclude any security for which a delisting occurs within the subsequent $\Delta t_{\text{out}}=5$ trading days. This procedure is safe because, under SEC rules, issuers must provide at least 10 days’ written notice to the exchange before filing Form 25 to delist securities [63].
@@
On the recent trading history, the closing auction must have executed on each day, and on each of those days, the daily traded volume must equal at least 1 % of shares outstanding and 1 % of market capitalization. This ensures that the stock is liquid enough to be traded with a low impact on the market. The day before the trade, we check that the number of shares outstanding is greater than 5M, to avoid micro-shares. Finally, on the day before trading, we require that each stock’s price lies between 10 \$ and 2,000 \$. Stocks trading at relatively low price levels incur high per-share transaction costs, whereas those at relatively high price levels often exhibit reduced trading activity, impeding efficient execution.
@@
Finally, to reduce redundancy, we keep only one class of shares for each company, the one with the largest market capitalization on the previous day. A similar filter is executed for those stocks with an in-sample correlation greater than 0.95 [65]. The final selection is done by ranking the stocks by their market capitalization and selecting the top $n=1,000$ stocks for every trading day $t$. The final panel, spanning January 1990 to December 2024, contains roughly 2.72.72.7 millions of stock–day pairs and approximately 4,0004,0004 , 000 unique stocks.
@@ 4.2 Training Procedure
The proposed architecture is designed to be independent of the number of assets $n$. This means that the same model can be applied to different subsets of assets without the need for retraining. However, we know from RMT that the filtering should depend on the aspect ratio $q=n/\Delta t_{\text{in}}$; therefore, during the training phase, the NN should be fed with examples of different $n$ sizes if we aim to obtain a model that generalizes well to larger portfolio sizes. This is done by randomly and uniformly sampling the number of assets $n$ from a [50,350][50,350][ 50 , 350 ] range in every batch of the training. The number of days $\Delta t_{\text{in}}$ is fixed to 120012001200 trading days; in fact, the presence of the module of Sec. 3.1 includes $2\Delta t_{\text{in}}$ trainable variables $(\alpha_{t},\beta_{t})$, which are associated with specific past lags. In the future version, this could be generalized to a variable number of past lags, but it would require a more complex architecture.
@@
According to the training procedure, the probability that NN will be trained on the same dataset is very low for $n$ sufficiently smaller than the investable universe on a given day. Consequently, continuously monitoring validation-set performance for early stopping becomes less critical. However, to further reduce the risk of overfitting, we used a decay factor of the learning rate, which is set to $10^{-4}$ and drops every batch by a factor of $0.99^{1/500}$, and we clip the gradients by norm to 1.01.01.0. The gradient is then optimized with the Adam optimizer with an automatic rescaling of the loss function. We trained all our models with 100100100 epochs, with 500500500 steps per epoch, and a batch size of 323232 samples. To ensure that the convergence is stable, for each of the 24 years, we trained our NN 10 times independently, so finally, we obtained 240 trained NNs.
@@
[…]
@@ 4.3.1 Lag-Transformed Returns
After the calibration, we can analyze the learned weighting factors $\alpha_{t}$ and $\beta_{t}$ of Eq. (14) as a function of the lag $t$ to understand how the NN is weighting and transforming the past returns. Given that we ran 101010 independent training runs for each of the 242424 years of the validations, we obtained 240240240 trained models. We can then compute the average of the weighting factors over the trained models, which are shown in Fig. 3.
@@
[…]
@@ 4.3.2 Eigenvalue Sensitivity Analysis
The counterintuitive result of the AO estimator [28, 27] is that a set of fixed eigenvalues can be used to obtain an estimator for the correlation matrix that outperforms all the RMT-based estimators for portfolio optimization. Moreover, such an estimator does not depend on the assets used in the calibration or the exchange, country, or time period. The authors of Ref. [28] hypothesized that the non-stationarity of the system is so strong that, unless eigenvectors are corrected, very little information can be extracted from the sample eigenvalues. In our case, we used an NN trained to attempt extracting specifically the non-stationary relationship between $\bm{C}_{\text{in}}$ and $\bm{C}_{\text{out}}$. If the above statement is true, and the NN does not overfit the calibration data, we should expect an output that is independent of the assets in the input.
[…]